Pack & Carry is debating whether to invest in new equipment to manufacture a line of high-quality luggage. The new equipment b. Compute the payback period for the investment in the new equipment to produce the new luggage line. (Round your answer to 1 decimal place.)

c. Compute the return on average investment for the investment in the new equipment to produce the new luggage line. (Round your percentage answer to 1 decimal place,(i.e., 0.1234 to be entered as 12.34.)

d.

Compute the total present value of the expected future annual cash inflows, discounted at an annual rate of 10 percent for the investment in the new equipment to produce the new luggage line. (Round your "PV factor" to 3 decimal places and final answer to the nearest dollar amount.)

e.

Compute the net present value of the proposed investment discounted at 10 percent for the investment in the new equipment to produce the new luggage line. (Round your "PV factor" to 3 decimal places and final answer to the nearest dollar amount.)
Answer #1

A. Cashflows per annum => Revenues - expenses actually paid. So, 967000 - 675000 = 292000.

B. For payback purpose we ignore Depreciation as it's a noncash item.

So, payback period = Initial Outlay/Cashflow per annum

= 900,000/292000 (from above)

= 3.08 Years or 3years, 1month.

C. ROA = Net income/initial outlay*100

Ie., 67000/900000*100 = 7.44%

D. Annual rate 10%. Useful life = 4 years

So PV of Cash Inflows => 292000*(PV factor for 4 years @ 10%) ie., similar 292000*3.170 = 925600.

E. So NPV of Project = PV OF Cash Inflows - initial capital outlay ie., 925600 - 900000 = +25600.